p(-6,-5), q(6,10)points.....the d=(p,q) and the midpoint thanks for ur time and help!
Answer by stanbon(75887)

p(-6,-5), q(6,10)points.....the d=(p,q) and the midpoint thanks for ur time and help!
-----
Are you saying (p,q) is the midpoint?
---
If so, p = (-6+6)/2 = 0
And q = (10-5)/2 = 2.5
----
(p,q) = (0,2.5)
========================
Cheers,
Stan H.